a cashier has 20 bills, all of which are $10 or $20 bills. The total value of the money is $330. How many of each type does the cashier have?
Found 2 solutions by richard1234, mananth:
Answer by richard1234(7193)

Let a,b be the number of $10 and $20 bills, respectively. We have the system of equations
10a + 20b = 330
a + b = 20 (since there are 20 bills)
We can multiply the second equation by 10 to obtain 10a + 10b = 200. Substituting into the first equation we have
(10a + 10b) + 10b = 330
200 + 10b = 330
b = 13, a = 7
So the cashier has seven $10 bills, 13 $20 bills.

Answer by mananth(16949)

$10 bill -------x
$20 bill -------y
...
x+y=20......................1
...
10+20y=330
/10
x+2y=33.....................2
multiply equation 1 by -1
-x-y-20
add this to (2)
-x+x-y+2y=-20+33
y = 13 $ 20 bills
x= 7 $10 bills
...
m.ananth@hotmail.ca